On June 30, 2021, Fly-By-Night Airlines leased a jumbo jet from Boeing Corporation. The terms of the lease require Fly-By-Night to make 20 annual payments of $400,000 on each June 30. Generally accepted accounting principles require this lease to be recorded as a liability for the present value of scheduled payments. Assume that a 7% interest rate properly reflects the time value of money in this situation. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.)

Required: 1. At what amount should Fly-By-Night record the lease liability on June 30, 2021, assuming that the first payment will be made on June 30, 2022?

table or calculator function-

payment-

n=

i=

PV/6-30-2021-

2. At what amount should Fly-By-Night record the lease liability on June 30, 2021, before any payments are made, assuming that the first payment will be made on June 30, 2021?

table or calc function-

payment-

n=

i=

PV-6/30/2021
Answer #1

Solution-

1)- n=20,i=7,

Lease liability to be recorded on June 30, 2021 = $400,000* Cumulative PV Factor at 7% for 20 periods

= $400,000* 10.59401 = $ 4,237,604

2)-n=20,i=7

Lease liability to be recorded on June 30, 2021 = $400,000* Cumulative PV Factor at 7% for 20 periods of annuity due

= $400,000* 11.33560 = $4,534,240